I'm not sure how to even set this up.
My son has $112 dollars made up of dimes, quarters and one dollar bills. The amount of one dollar bills is 52 dollars and the amount of quarters is two times the amount of dimes. How many does he have of each?
I got as far as 112-52= 70 (to be left with only the coins to resolve), but I didn't understand how to get any further. Please help!
Thanks!

Answer by stanbon(75887)

My son has $112 dollars made up of dimes, quarters and one dollar bills. The amount of one dollar bills is 52 dollars and the amount of quarters is two times the amount of dimes. How many does he have of each?
--------------------------------------------------
Value Equation: 10d + 25q + 5200 = 11200 cents
10d + 25q = 6000 cents
q = 2d
10d + 50d = 6000
60d = 6000
d = 100 (# of dimes)
q = 2d = 200 (# of quarters)
b = 52 (# of dollars)
========================================
Cheers,
Stan H.